OpenStudy (anonymous):

find the value of Q in the following system so that the solution to the system is {(x,y) : x-3y=4} x-3y=4 2x+Qy=8

OpenStudy (freckles):

you want 2x+Qy=8 to be the same as x-3y=4

OpenStudy (freckles):

hint: try dividing 2x+Qy=8 by 2